# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ 2.3 Maximum Property of the Binary Proximal Outcome
If, however, $Y_{t,\Delta}$ is defined by whether an event is sustained for a period of time, then (3) may not hold. For example, in the HeartSteps II MRT, the binary proximal outcome is whether an activity bout occurred during the 30-minute window. Consider an activity bout to be a period of sustained activity of 90 seconds or longer. Suppose an individual started activity 45 seconds prior to decision point $t+1$, and concluded activity 45 seconds after $t+1$. In this case, their $Y_{t,\Delta=2}$ would be 1, but there is no straightforward definition of $R_{t+1}$ to satisfy (3). In fact, there is no activity bout of 90 seconds that is completely inside either of the sub-intervals $(t,t+1)$ and $(t+1,t+2)$. To generalize our method to such cases, we can introduce a double-subscript version of the sub-outcome $R$ along with a modified maximum property to determine the first decision point, within the proximal outcome window, at which the event has occurred. However, this generalization would make the notation too cumbersome without adding substantial novelty. Therefore, we focus on the current version of maximum property (3), and present how to modify the sub-outcomes and the maximum property to generalize our method in Supplementary Material A.
@@ 2.4 Identification of Parameters
Assumption 1 is standard for causal inference settings without interference. It may be violated if the mobile health intervention incorporates community features or the participants are clustered, in which case a causal inference framework that allows interference should be used instead [28]. (There is no community feature in the Drink Less app, our motivating example.) Assumption 2 is always valid for MRTs because the randomization probability of $A_{t}$ depends (at most) on $H_{t}$ by study design. Assumption 3 is valid as long as the researchers do not set the randomization probability to exactly 0 or 1 when $I_{t}=1$. Assumption 4 is the potential outcomes version of the maximum property (3).
Let $\mathbbm{1}(\cdot)$ denote the indicator function, and define $10^{0}=$ 1. In the Supplementary Material C, we prove that under the above 4 assumptions the casual excursion effect in (2) can be expressed using observed data distribution:
@@
We define $\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}=1$ if $\Delta=1$. Equation (4) will be used to establish the asymptotic property of our estimator. We will slightly abuse the notation and denote the right hand side of (4) by $\text{CEE}(S_{t})$.
3 An Estimator Using Per-Decision Inverse Probability Weights
@@
Our estimator for $\beta$, denoted by $\hat{\beta}$, requires a working model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$, which characterizes the expected outcome under no treatment between $t$ and $t+\Delta$ given the history $H_{t}$ and $I_{t}=1$. We use $\exp\{g(H_{t})^{T}\alpha\}$ as the working model, where $g(H_{t})$ is a vector of features constructed from $H_{t}$ and $\alpha$ is a $q$-dimensional nuisance parameter. As will be established in Theorem 1, $\hat{\beta}$ is consistent and asymptotically normal regardless of whether the working model $\exp\{g(H_{t})^{T}\alpha\}$ is a correct model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$.
Our estimator $\hat{\beta}$, along with the nuisance estimator $\hat{\alpha}$, is obtained by solving the estimating equation $\frac{1}{n}\sum_{i=1}^{n}U_{i}(\alpha,\beta)=0$, with the estimating function $U_{i}(\alpha,\beta)$ defined as
@@
The estimating function is multiplied by $I_{it}$ because the causal excursion effect only concerns the decision points where individuals are eligible to be randomized. The factor $e^{-A_{it}S_{it}^{T}\beta}$ is analogous to the blipping-down technique in multiplicative structural nested mean models [23]. The error term $\epsilon_{it}(\alpha,\beta)$ is defined as $\epsilon_{it}(\alpha,\beta)=Y_{it,\Delta}-e^{g(H_{it})^{T}\alpha+A_{it}S_{it}^{T}\beta}$, where $Y_{it,\Delta}$ is the proximal outcome $Y_{t,\Delta}$ for the $i$-th individual. The $M_{it}$ is defined as $M_{it}=\left\{\frac{\tilde{p}_{t}(S_{it})}{p_{t}(H_{it})}\right\}^{A_{it}}\left\{\frac{1-\tilde{p}_{t}(S_{it})}{1-p_{t}(H_{it})}\right\}^{1-A_{it}}$, which is similar to the stabilized IPW [24] and is employed by other estimators for causal excursion effects [18, 11]. Here $\tilde{p}_{t}(S_{it})\in(0,1)$ is similar to the numerator of a stabilized weight and can be chosen to improve efficiency by making $\tilde{p}_{t}(S_{it})$ as close to $p_{t}(H_{it})$ as possible. For example, if $p_{t}(H_{it})$ is a constant or depends only on $S_{it}$, one can simply set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ which makes $M_{it}\equiv 1$; if $p_{t}(H_{it})$ depends on variables in $H_{it}\setminus S_{it}$, one should not set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ but can set $\tilde{p}_{t}(S_{it})$ to be the prediction from a logistic regression with $A_{it}$ being the response and $S_{it}$ being the predictor. The $W_{it}$ is defined as $W_{it}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{i,t+s}=0)}$ which is the per-decision inverse probability weight that we will elaborate on later.
@@
Assume the causal effect model (5) and Assumptions 1-4 in Section 2.4 hold. Let $\beta^{*}$ denote the true value of $\beta$ corresponding to the data generating distribution $P_{0}$. Define $\dot{U}_{i}(\alpha,\beta)=\partial U_{i}(\alpha,\beta)/\partial(\alpha^{T},\beta^{T})$. Under regularity conditions, $\hat{\beta}$ is consistent for $\beta^{*}$ and $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with mean zero and variance-covariance matrix $\Sigma$, and $\Sigma$ can be consistently estimated by the lower block diagonal $(p\times p)$ entry of the matrix
$\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1}\left\{\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1,T}.$
@@
For the estimand $\beta^{*}$, [11] proposed an estimator called the estimator for marginal excursion effect (EMEE). The EMEE is the same as our estimator except that $W_{it}$ is replaced by $W_{it}^{\prime}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}\right\}$. The $W_{it}^{\prime}$ is a product of $(\Delta-1)$ inverse probability weights (IPW) to account for the fact that the causal excursion effect (2) concerns excursions under treatment trajectory $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$. Let us examine how the EMEE weighs the $(i,t)$-decision point by $W_{it}^{\prime}$ and later contrast with our estimator. Case 1: suppose for some $t+1\leq j\leq t+\Delta-1$, $A_{ij}=1$ (Table 1, Examples 1a and 1b). Because this is different from the treatment trajectory for the excursion, the $(i,t)$-decision point is assigned weight $W_{it}^{\prime}=0$ by EMEE. Case 2: suppose $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$ (Table 1, Examples 2a and 2b). In this case $W_{it}^{\prime}>1$ and the $(i,t)$-decision point is upweighted to account for itself and similar decision points that fall into Case 1 and thus have weight 0.
@@
The key innovation in our proposed estimator is $W_{it}$, which improves efficiency over EMEE. We call $W_{it}$ the per-decision inverse probability weight because the inclusion of each of the $(\Delta-1)$ IPW terms is determined by each decision point: for each $j$ from $t+1$ to $t+\Delta-1$, if $\max_{1\leq s\leq j-t}R_{i,t+s}=0$ then $\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}$ is included in $W_{it}$; otherwise it is not included. Therefore, $W_{it}$ includes in its product the IPW terms from $j=t+1$ to right before the first occurrence of some $R_{j^{*}}=1$. By not including the IPW terms after some $R_{j^{*}}=1$, the per-decision IPW does not account for the treatment assignments after $j^{*}$. This is appropriate because the maximum property (3) implies that the treatment assignments after $j^{*}$ doesn’t affect the value of $Y_{t,\Delta}$ (which has to equal 1). To see how our estimator improves efficiency over the EMEE, let us now examine how our estimator weighs the $(i,t)$-decision point by $W_{it}$. For Case 1 described above, if some $R_{j}=1$ occurs before the first $A_{k}=1$ (Table 1, Example 1b), the $(i,t)$-decision point will have a positive weight. For Case 2 described above, if the first $R_{j^{*}}=1$ occurs for $j^{*}<t+\Delta-1$ (Table 1, Example 2b), then $W_{it}$ will include fewer terms than $W_{it}^{\prime}$ and thus be smaller, because $W_{it}=\prod_{j=t+1}^{j^{*}}\frac{1}{1-p_{j}\left(H_{ij}\right)}<\prod_{j=t+1}^{t+\Delta-1}\frac{1}{1-p_{j}\left(H_{ij}\right)}=W_{it}^{\prime}$.
@@
To summarize, by capitalizing on the maximum property (3), the per-decision IP weight $W_{it}$ makes use of some decision points that would be otherwise discarded because $W_{it}^{\prime}=0$ (Example 1b) and reduces the weight of some decision points (Example 2b), and this leads to the efficiency improvement. The efficiency improvement of our estimator is larger when a larger proportion of the decision points belong to the above two types, which happens when the proximal outcome window length $\Delta$ is large, when the randomization probability is large, or when the baseline success probability (i.e., the success probability of the proximal outcome under no treatment) is larger. When the proximal outcome window is shorter than the distance between two decision points so that $\Delta=1$, our estimator becomes equivalent to EMEE. In Supplementary Material E we provided a theoretical analysis for the relative efficiency between the proposed estimator and the EMEE under simplifying assumptions.
@@
Related techniques of reducing IP weights to improve efficiency were developed for other settings. In off-policy reinforcement learning, the per-decision importance sampling technique was proposed to construct IP weights related to ours by leveraging the fact that earlier outcomes cannot depend on later treatments [14]. In Cox marginal structural model, to determine a subject’s contribution to a risk-set at time $t$, a stabilized weight was proposed which involves IP weights not through the end of the follow-up but only up till time $t$ [24]. Those settings are analogous to a special case in our setting ($T=1$), i.e., when focusing on a single proximal outcome that decomposes into a sequence of sub-outcomes. Our contribution is to extend the per-decision IPW technique to estimating causal excursion effects on longitudinal proximal outcomes. In Supplementary Material B, we provide a detailed explication on how our method connects to and differs from the literature.
@@
[…]

### v3

@@ 2.3 Maximum Property of the Binary Proximal Outcome
If, however, $Y_{t,\Delta}$ is based on a cumulative event, then (3) may not hold. For example, in the HeartSteps I MRT, the binary proximal outcome is whether a participant takes more than, say, 10,000 steps in the next 24 hours (a proximal outcome window spanning over $\Delta=5$ decision points). In this case, no sub-outcomes $R_{t+1}$ can satisfy (3) because the total step count is cumulative instead of instantaneous. We generalize the maximum property to accommodate such settings in Supplementary Material B by introducing a double-subscript version of the sub-outcome $R$. However, this generalization would make the notation too cumbersome without adding substantial novelty, so we focus on the current version of the maximum property (3) in the sequel when we introduce the methods.
@@ 2.4 Identification of Parameters
Assumption 1 is standard for causal inference settings without interference. It may be violated if the mobile health intervention incorporates community features or the participants are clustered, in which case a causal inference framework that allows interference should be used instead (Hudgens and Halloran, 2008; Shi and others, 2023). (There is no community feature in the Drink Less app, our motivating example.) Assumption 2 is always valid for MRTs because the randomization probability of $A_{t}$ depends (at most) on $H_{t}$ by study design. Assumption 3 is valid as long as the researchers do not set the randomization probability to exactly 0 or 1 when $I_{t}=1$. Assumption 4 is the potential outcomes version of the maximum property (3).
Let $\mathbbm{1}(\cdot)$ denote the indicator function, and define $10^{0}=$ 1. In Supplementary Material C, we prove that under the above 4 assumptions the casual excursion effect in (2) can be expressed using observed data distribution:
@@
We define $\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}=1$ if $\Delta=1$. Equation (4) will be used to establish the asymptotic property of our estimator. We will slightly abuse the notation and denote the right-hand side of (4) by $\text{CEE}(S_{t})$.
3 pd-EMEE: An Estimator Using Per-Decision Inverse Probability Weights
@@
Our first estimator for $\beta$ is denoted by $\hat{\beta}$ and referred to as pd-EMEE (a name we will explain later). It requires a working model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$, which characterizes the expected outcome under no treatment between $t$ and $t+\Delta$ given the history $H_{t}$ and $I_{t}=1$. We use $\exp\{g(H_{t})^{T}\alpha\}$ as the working model, where $g(H_{t})$ is a vector of features constructed from $H_{t}$ and $\alpha$ is a $q$-dimensional nuisance parameter. As will be established in Theorem 1, $\hat{\beta}$ is consistent and asymptotically normal regardless of whether the working model $\exp\{g(H_{t})^{T}\alpha\}$ is a correct model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$.
The estimator $\hat{\beta}$, along with the nuisance estimator $\hat{\alpha}$, is obtained by solving an estimating equation $\frac{1}{n}\sum_{i=1}^{n}U_{i}(\alpha,\beta)=0$, with the estimating function $U_{i}(\alpha,\beta)$ defined as
@@
The estimating function is multiplied by $I_{it}$ because the causal excursion effect only concerns the decision points where individuals are available to be randomized. The factor $e^{-A_{it}S_{it}^{T}\beta}$ is analogous to the blipping-down term in multiplicative structural nested mean models (Robins, 1994). The error term $\epsilon_{it}(\alpha,\beta)$ is defined as $\epsilon_{it}(\alpha,\beta)=Y_{it,\Delta}-e^{g(H_{it})^{T}\alpha+A_{it}S_{it}^{T}\beta}$, where $Y_{it,\Delta}$ is the proximal outcome $Y_{t,\Delta}$ for the $i$-th individual. The $M_{it}$ is defined as $M_{it}=\left\{\frac{\tilde{p}_{t}(S_{it})}{p_{t}(H_{it})}\right\}^{A_{it}}\left\{\frac{1-\tilde{p}_{t}(S_{it})}{1-p_{t}(H_{it})}\right\}^{1-A_{it}}$, which is similar to the stabilized IPW (Robins and others, 2000) and is employed by other estimators for causal excursion effects (Boruvka and others, 2018; Qian and others, 2021). Here $\tilde{p}_{t}(S_{it})\in(0,1)$ is similar to the numerator of a stabilized weight and can be chosen to improve efficiency by making $\tilde{p}_{t}(S_{it})$ as close to $p_{t}(H_{it})$ as possible. For example, if $p_{t}(H_{it})$ is a constant or depends only on $S_{it}$, one can simply set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ which makes $M_{it}\equiv 1$; if $p_{t}(H_{it})$ depends on variables in $H_{it}\setminus S_{it}$, one should not set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ but can set $\tilde{p}_{t}(S_{it})$ to be the prediction from a logistic regression with $A_{it}$ being the response and $S_{it}$ being the predictor. The $W_{it}$ is defined as $W_{it}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{i,t+s}=0)}$ which is the per-decision inverse probability weight that we will elaborate on later.
@@
Assume the causal effect model (5) and Assumptions 1-4 in Section 2.4 hold. Let $\beta^{*}$ denote the true value of $\beta$ corresponding to the data generating distribution $P_{0}$. Define $\partial_{\beta}U_{i}(\alpha,\beta)=\partial U_{i}(\alpha,\beta)/\partial(\alpha^{T},\beta^{T})$. Under regularity conditions, $\hat{\beta}$ is consistent for $\beta^{*}$ and $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with mean zero and variance-covariance matrix $\Sigma$, and $\Sigma$ can be consistently estimated by the lower block diagonal $(p\times p)$ entry of the matrix
$\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}U_{i}(\hat{\alpha},\hat{\beta})\bigg{\}}^{-1}\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\bigg{\}}\bigg{\{}\big{[}\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}U_{i}(\hat{\alpha},\hat{\beta})\big{]}^{-1}\bigg{\}}^{T}.$
@@
For the estimand $\beta^{*}$, Qian and others (2021) proposed an estimator called the estimator for marginal excursion effect (EMEE). The EMEE is the same as our estimator except that $W_{it}$ is replaced by $W_{it}^{\prime}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}\right\}$. The $W_{it}^{\prime}$ is a product of $(\Delta-1)$ inverse probability weights (IPW) to account for the fact that the causal excursion effect (2) concerns excursions under treatment trajectory $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$. Let us examine how the EMEE weighs the $(i,t)$-decision point by $W_{it}^{\prime}$ and later contrast with our estimator. Case 1: suppose for some $t+1\leq j\leq t+\Delta-1$, $A_{ij}=1$ (Table 4, Examples 1a and 1b). Because this is different from the treatment trajectory for the excursion, the $(i,t)$-decision point is assigned weight $W_{it}^{\prime}=0$ by EMEE. Case 2: suppose $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$ (Table 4, Examples 2a and 2b). In this case, $W_{it}^{\prime}>1$ and the $(i,t)$-decision point is upweighted to account for itself and similar decision points that fall into Case 1 and thus have weight 0.
@@
The key innovation in our proposed estimator is $W_{it}$, which improves efficiency over EMEE. We call $W_{it}$ the per-decision inverse probability weight, and $\hat{\beta}$ the pd-EMEE method, because the inclusion of each of the $(\Delta-1)$ IPW terms is determined by each decision point: for each $j$ from $t+1$ to $t+\Delta-1$, if $\max_{1\leq s\leq j-t}R_{i,t+s}=0$ then $\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}$ is included in $W_{it}$; otherwise it is not included. Therefore, $W_{it}$ includes in its product the IPW terms from $j=t+1$ to right before the first occurrence of some $R_{j^{*}}=1$. By not including the IPW terms after this $R_{j^{*}}=1$, the per-decision IPW does not account for the treatment assignments after $j^{*}$. This is appropriate because the maximum property (3) implies that the treatment assignments after $j^{*}$ don’t affect the value of $Y_{t,\Delta}$ (which has to equal 1). To see how our estimator improves efficiency over the EMEE, let us now examine how our estimator weighs the $(i,t)$-decision point by $W_{it}$. For Case 1 described above, if some $R_{j}=1$ occurs before the first $A_{k}=1$ (Table 4, Example 1b), the $(i,t)$-decision point will have a positive weight. For Case 2 described above, if the first $R_{j^{*}}=1$ occurs for $j^{*}<t+\Delta-1$ (Table 4, Example 2b), then $W_{it}$ will include fewer terms than $W_{it}^{\prime}$ and thus be smaller, because $W_{it}=\prod_{j=t+1}^{j^{*}}\frac{1}{1-p_{j}\left(H_{ij}\right)}<\prod_{j=t+1}^{t+\Delta-1}\frac{1}{1-p_{j}\left(H_{ij}\right)}=W_{it}^{\prime}$.
@@
To summarize, by capitalizing on the maximum property (3), the per-decision IP weight $W_{it}$ makes use of some decision points that would be otherwise discarded because $W_{it}^{\prime}=0$ (such as Example 1b) and reduces the weight of some other decision points (such as Example 2b), and this leads to the efficiency improvement. The efficiency improvement of $\hat{\beta}$ over EMEE is larger when a larger proportion of the decision points belong to the above two types, which happens when the proximal outcome window length $\Delta$ is large, when the randomization probability is large, or when the baseline success probability (i.e., the success probability of the proximal outcome under no treatment) is large. When the proximal outcome window is shorter than the distance between two decision points so that $\Delta=1$, $\hat{\beta}$ is equivalent to EMEE. We will use “relative efficiency” to quantify the efficiency improvement, with relative efficiency between two estimators $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$ defined as $\text{Var}(\hat{\theta}_{2})/\text{Var}(\hat{\theta}_{1})$. In Supplementary Material E.2 we provided a theoretical analysis for the relative efficiency between the proposed estimator and the EMEE under simplifying assumptions.
@@
Related techniques of reducing IP weights to improve efficiency were developed for other settings. In off-policy reinforcement learning, the per-decision importance sampling technique was proposed to construct IP weights related to ours by leveraging the fact that earlier outcomes cannot depend on later treatments (Precup, 2000). In Cox marginal structural model, to determine a subject’s contribution to a risk-set at time $t$, a stabilized weight was proposed which involves IP weights not through the end of the follow-up but only up till time $t$ (Robins and others, 2000). Those settings are analogous to a special case in our setting ($T=1$), i.e., when focusing on a single proximal outcome that decomposes into a sequence of sub-outcomes. Our contribution is to extend the per-decision IPW technique to estimate causal excursion effects on longitudinal proximal outcomes. In Supplementary Material F, we provide a detailed explication on how our method connects to and differs from the literature.
@@
(Robustness against a misspecified working model.) Consistency of $\hat{\beta}$ only requires that $S_{t}^{T}\beta$ be a correct model for the causal excursion effect $\text{CEE}(S_{t})$. It does not require a correct working model $g(H_{t})^{T}\alpha$; i.e., it does not require $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}=\exp\{g(H_{t})^{T}\alpha\}$ to hold for some $\alpha$. This robustness property is desired because it is virtually impossible to correctly specify a model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$ when the number of decision points for an MRT is moderate to large. Different choices of $g(H_{t})$ affect the variance of $\hat{\beta}$. See Section 4 and Remark 3 on how the projection idea leads to a more efficient choice of $g(H_{t})^{T}\alpha$.